OpenStudy (anonymous):

3^5 sqrt (x+2)^3 + 3 = 27

OpenStudy (anonymous):

\[3^5 \sqrt{ (x+2)^3} + 3 = 27\]Is this the equation?

OpenStudy (anonymous):

\[3^5 \sqrt{ (x+2)^3 + 3} = 27\]Or this?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Anyway you could help would be appreciated!

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[3^5 \sqrt{ (x+2)^3} + 3 = 27\]\[3^5 \sqrt{ (x+2)^3} + 3 \color{red}{-3}= 27 \color{red}{-3}\]\[3^5 \sqrt{ (x+2)^3} = 24\]\[\color{red}{ \color{black}{3^5 \sqrt{ (x+2)^3} }\over 3^5} = {\color{red}{\color{black}{24} \over 3^5}}\]\[ \sqrt{ (x+2)^3} = {24 \over 3^5}\]\[\color{red}{(}\sqrt{ (x+2)^3}\color{red}{)^2} = \color{red}{(}{24 \over 3^5}\color{red}{)^2}\]\[(x+2)^3=\frac{24^2}{3^{10}}\]\[\color{red}((x+2)^3\color{red}{)^{1/3}}=\color{red}(\frac{24^2}{3^{10}}\color{red}{)^{1/3}}\]\[x+2=\frac{24^{2/3}}{3^{10/3}}\]\[x+2\color{red}{-2}=\frac{24^{2/3}}{3^{10/3}}\color{red}{-2}\]\[x=\frac{24^{2/3}}{3^{10/3}}-2\]I'll leave the calculator stuff to you